# Automatic Search of Multiword Place Names on Historical Maps

### v1

@@ 1. Introduction
Previous work on document layout understanding aims to link multiple words into phrases based on spatial relations of detected bounding boxes (Binmakhashen and Mahmoud, 2019). However, existing work mainly focuses on documents such as receipts or reports comprising horizontally/vertically aligned text labels. In the case of scanned historical maps, alignments of text labels are not uniform within scanned maps and across different scanned maps. Figure  LABEL:fig:linking_challenges shows three main challenges in linking text labels of place phrases on historical maps: 1) the density of text labels on maps, 2) the inconsistency with which multiword place names are laid out, and 3) the presence of unrelated text labels appearing between text labels that are in multiword phrases.
To overcome these challenges, we propose a simple and efficient approach for searching multiword place names among recognized text from historical maps. Figure  1 illustrates this approach, which consists of two steps: 1) constructing a minimum spanning tree (MST), which links text labels into potential phrases, and 2) searching for the multiword place name in this tree.
@@ 2.1. Linking Multiword Place Names
To understand how likely pairs of text labels are to be adjacent in a multiword phrase, we examine explicit visual features of text labels, including distance, height (i.e., font size), orientation, and capitalization. In the most straightforward cases, text labels within a multiword place name are closer to each other than text labels that belong to other place names. Thus, we leverage the distance between text labels as an important feature. However, there are challenging cases where unrelated text labels appear between text labels that form a multiword place name. In such cases, human readers can reason that these text labels are not part of the name because they might have a different font size, angle of orientation, or capitalization to the text labels in the multiword place name. For example, in Figure LABEL:fig:linking_challenges, the unrelated text labels appearing between the multiword name “QUEENS COUNTY” have a smaller font size and are not in all capitals, and the text label between the multiword name “Gt. Barrier Reef” is oriented at an angle nearly perpendicular to the text labels in that multiword name. From these observations, we define a heuristic called an edge cost function, which represents how unlikely two text labels $l_{i}$ and $l_{j}$ are to be adjacent in a multiword phrase. This edge cost function is based on four explicit visual features of the text labels. The first is the minimum spatial distance between the bounding boxes of the two text labels on the map image, denoted as $d(l_{i},l_{j})$. The second is the ratio of the height of the label with the taller bounding box to that of the label with the shorter bounding box, minus $1$. This feature, denoted as $h(l_{i},l_{j})$, will be minimized at 00 for labels with the same height and, by extension, similar font sizes. The third feature, denoted as $a(l_{i},l_{j})$, is the sine of an angle difference between the axes of two bounding boxes, which will range from 00 to $1$ and be minimized when the axes of the bounding boxes of two labels are aligned. The final feature, denoted as $c(l_{i},l_{j})$, represents the difference in capitalization between the labels. The value of $c(l_{i},l_{j})$ is 00 when either the two labels are all uppercase, or neither is all uppercase, or $1$ if one is all uppercase and the other is not.
@@ 15 Most
Table  2 shows statistics of the time-sequenced map query results. These queries retrieved an average of 80.22 maps matching multiword place names per query, and these maps spanned an average time range of 205 years. Figure  3 shows maps sampled from the 42 time-sequenced map results of a query for “Sault Ste. Marie”. These results show that this approach can be used to retrieve many historical maps of places that use multiword names, revealing changes in each place across wide ranges of history.

### v2

@@ 1. Introduction
Figure 1. Challenges of linking together multiword place names on historical maps.
@@
Previous work on document layout understanding aims to link multiple words into phrases based on spatial relations of detected bounding boxes (Binmakhashen and Mahmoud, 2019). However, existing work mainly focuses on documents such as receipts or reports comprising horizontally/vertically aligned text labels. In the case of scanned historical maps, alignments of text labels are not uniform within scanned maps and across different scanned maps. Figure  1 shows three main challenges in linking text labels of place phrases on historical maps: 1) the density of text labels on maps, 2) the inconsistency with which multiword place names are laid out, and 3) the presence of unrelated text labels appearing between text labels that are in multiword phrases.
To overcome these challenges, we propose a simple and efficient approach for searching multiword place names among recognized text from historical maps. Figure  2 illustrates this approach, which consists of two steps: 1) constructing a minimum spanning tree (MST), which links text labels into potential phrases, and 2) searching for the multiword place name in this tree.
@@ 2.1. Linking Multiword Place Names
To understand how likely pairs of text labels are to be adjacent in a multiword phrase, we examine explicit visual features of text labels, including distance, height (i.e., font size), orientation, and capitalization. In the most straightforward cases, text labels within a multiword place name are closer to each other than text labels that belong to other place names. Thus, we leverage the distance between text labels as an important feature. However, there are challenging cases where unrelated text labels appear between text labels that form a multiword place name. In such cases, human readers can reason that these text labels are not part of the name because they might have a different font size, angle of orientation, or capitalization to the text labels in the multiword place name. For example, in Figure 1, the unrelated text labels appearing between the multiword name “QUEENS COUNTY” have a smaller font size and are not in all capitals, and the text label between the multiword name “Gt. Barrier Reef” is oriented at an angle nearly perpendicular to the text labels in that multiword name. From these observations, we define a heuristic called an edge cost function, which represents how unlikely two text labels $l_{i}$ and $l_{j}$ are to be adjacent in a multiword phrase. This edge cost function is based on four explicit visual features of the text labels. The first is the minimum spatial distance between the bounding boxes of the two text labels on the map image, denoted as $d(l_{i},l_{j})$. The second is the ratio of the height of the label with the taller bounding box to that of the label with the shorter bounding box, minus $1$. This feature, denoted as $h(l_{i},l_{j})$, will be minimized at 00 for labels with the same height and, by extension, similar font sizes. The third feature, denoted as $a(l_{i},l_{j})$, is the sine of an angle difference between the axes of two bounding boxes, which will range from 00 to $1$ and be minimized when the axes of the bounding boxes of two labels are aligned. The final feature, denoted as $c(l_{i},l_{j})$, represents the difference in capitalization between the labels. The value of $c(l_{i},l_{j})$ is 00 when either the two labels are all uppercase, or neither is all uppercase, or $1$ if one is all uppercase and the other is not.
@@ 15 Most
Table  2 shows statistics of the time-sequenced map query results. These queries retrieved an average of 80.22 maps matching multiword place names per query, and these maps spanned an average time range of 205 years. Figure  4 shows maps sampled from the 42 time-sequenced map results of a query for “Sault Ste. Marie”. These results show that this approach can be used to retrieve many historical maps of places that use multiword names, revealing changes in each place across wide ranges of history.